OpenStudy (anonymous):

Please help - Monthly payments on a 30 year basis for a $32,000 loan at 9% amount to monthly payments of $257.48. Prof gave me formula: m= (Rate*Cost)/(1200*(1-(1+((rate/1200)^(-12* years))))) I keep plugging it in but I am NOT getting what he got?!? Can someone explain how this formula works?

OpenStudy (anonymous):

let me type it in with judicious use of parentheses and see what we get hold on

OpenStudy (anonymous):

ok i think the formula has a mistake in it

OpenStudy (anonymous):

where did the 1200 come from?

OpenStudy (anonymous):

My professor gave it to me. I have no idea.....I have tried solving it every which way but I do not get his answer.

OpenStudy (anonymous):

the one i know is \[\frac{3200\times \frac{.09}{12}}{1-(1+\frac{.09}{12})^{-360}}\]

OpenStudy (anonymous):

make sure to use -360 and not -12 *30 because that is liable to mess you up in the calculator

OpenStudy (anonymous):

ok (i guess maybe it was a really bad typo?) Thank you so much!

OpenStudy (anonymous):

here is what i got, it matches up as you can see http://www.wolframalpha.com/input/?i=32000%28.09%2F12%29%2F%281-%281%2B%28.09%2F12%29%29^%28-360%29%29%29

OpenStudy (anonymous):

Yeah! Ok, I'll shoot him an email to let him know and continue forward this way. Thanks again!!!!!!